OpenStudy (anonymous):

what are roots of the polynomial F(x)=3x^2-13x+4

OpenStudy (baldymcgee6):

do you know how to factor it?

OpenStudy (anonymous):

yea i got the answer now

OpenStudy (anonymous):

@chellyd123 it is simple let f(x)=0 so 3x^2-13x+4=0 3x^2-12x-x+4=0 3x(x-4)-1(x-4)=0 (3x-1)(x-4)=0 3x-1=0 or x-4=0 x=1/3 or x=4 roots are 1/3 and 4